OpenStudy (lyssburleson):

Translate this sentence into an equation. The product of Hector's score and 2 is 30. Use the variable h to represent Hector's score.

OpenStudy (vuriffy):

Do you know what product means?

OpenStudy (vuriffy):

Also, do you know what sum is?

OpenStudy (vuriffy):

Sum is addition. Product is multiplication So, (h*2) is means = Can you configure it now altogether?

OpenStudy (lyssburleson):

So i just write out the equation then? which would be h*2=30?

OpenStudy (vuriffy):

Correct. h*2 = 30.

OpenStudy (lyssburleson):

Thank you!

OpenStudy (vuriffy):

Of course, any time.

OpenStudy (lyssburleson):

Could you help me with another one? @Vuriffy

OpenStudy (lyssburleson):

The question is: 66 decreased by twice Carlos's age. I'm thinking it's 66-2c? @Vuriffy

OpenStudy (vuriffy):

That is correct.

OpenStudy (vuriffy):

Say, C = 2. 66 - 2(2) = D 66 - 4 = 62. Works quite well. Decreases by twice Carlo's age.

OpenStudy (lyssburleson):

thank you again :)